OpenStudy (anonymous):

Solve this system of equations using any method. 4x – y = 5 3x + y = 16 (3, 7) (3, 5) (4, 11) (4, 4)

OpenStudy (anonymous):

On the second equation, what does y equal?

OpenStudy (anonymous):

its (3,7)

OpenStudy (anonymous):

@Christhewiz187 We try to help people, not just give them the answers, when they take tests the answers don't just appear.

OpenStudy (anonymous):

ya they ask us then they do appear

OpenStudy (anonymous):

@satellite73 Am I right?

OpenStudy (marigirl):

4x – y = 5 and 3x + y = 16 Lets do substitution (only because I like it better lol) let us take the second equation and make y the subject. once you have done this replace the y (i.e substitute) in the first equation by this expression. Then u will see that you only have x as the only variable in your equation which will allow you to solve for x and find the x intercept where both the graphs cross over.

OpenStudy (anonymous):

I asked to learn how to do it but thanks

OpenStudy (anonymous):

'Preciate it

OpenStudy (anonymous):

@Spiffy2wice Can you answer my first question? I also like substitution @marigirl :D What does y equal in the second equation?

OpenStudy (anonymous):

she didn't specify so im not sure

OpenStudy (marigirl):

from your second equation, 3x + y = 16 can u make y the subject. (i.e y on the left side by it self)